OpenStudy (anonymous):

What is the distance between P(-4,3) and Q(6,1). Round tot he nearest tenth

OpenStudy (zzr0ck3r):

for (a,b) and (c,d) the distance is \(D=\sqrt{(a-c)^2+(b-d)^2}\)

OpenStudy (anonymous):

so what do i puyt in place of the letters in the formula

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

well instead of (a,b) and (c,d) you have (-4,3) and (6,1)

OpenStudy (zzr0ck3r):

so a=-4 b=3 c=6 d=1

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

hold on a secx thoe k

OpenStudy (anonymous):

so subtract and then divide

OpenStudy (zzr0ck3r):

what? there is no divide

OpenStudy (anonymous):

ok so sybtract nd add

OpenStudy (anonymous):

-4 - 3=-7 an 6-1=5

OpenStudy (anonymous):

so do i add -7+5

OpenStudy (anonymous):

answer would be -12 right

OpenStudy (zzr0ck3r):

where did you get -7?

OpenStudy (zzr0ck3r):

i didnt say (a-b) I said (a-c)

OpenStudy (anonymous):

ohh lol my bad i mean -4 -3=-1

OpenStudy (zzr0ck3r):

\(\sqrt{(-4-6)^2+(3-1)^2}=\sqrt{(-10)^2+2^2}= \sqrt{100+4}=\sqrt{104}\)

OpenStudy (anonymous):

umm u sure thats right because i did it a didffrent way

OpenStudy (zzr0ck3r):

(a-c) = (-4-6) = -10

OpenStudy (anonymous):

so for the first one is -10

OpenStudy (anonymous):

so i add that to what

OpenStudy (anonymous):

the lasy part

OpenStudy (zzr0ck3r):

\(\sqrt{(-4-6)^2+(3-1)^2}=\sqrt{(-10)^2+2^2}= \sqrt{100+4}=\sqrt{104}\approx10.2\)

OpenStudy (zzr0ck3r):

specifically what part?

OpenStudy (anonymous):

squared 100+4= squareed 104

OpenStudy (zzr0ck3r):

what is 100+4?

OpenStudy (anonymous):

thas what u put squared 100+4

OpenStudy (zzr0ck3r):

\(\sqrt{100+4}=\sqrt{104}\)

OpenStudy (anonymous):

yea that

OpenStudy (zzr0ck3r):

all I did was add 4 to 100

OpenStudy (zzr0ck3r):

I think you might need to go back a few chapters and catch up.

OpenStudy (anonymous):

ohhh 10.2 is the answer

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Are these the steps: 1. Take the Absolute Difference between the two X points, -4 and 6. The Absolute Difference is 10. 2. Square that number. 10 Squared is 100 3. Take the Absolute Difference between the two Y points 3 and 1. The absolute Difference is 2. 4. Square that number. 2 Squared is 4. 5. Add the two. 100 plus 4 is 104 6. take the Square Root of 104. The final answer is 10.1980390272 Rounded to the nearest Tenth. 10.2

OpenStudy (zzr0ck3r):

\(\sqrt{104}=\sqrt{4*26}=\sqrt{4}\sqrt{26}=2\sqrt{26}\approx 10.2\)

OpenStudy (anonymous):

yea

OpenStudy (zzr0ck3r):

you are correct

OpenStudy (anonymous):

ok thankss

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

XD